# Understanding Fairness in Software Engineering: Insights from Stack Exchange Sites

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: 18th ACM/IEEE International Symposium on Empirical Software Engineering and Measurement; October 20–25, 2024; Barcelona, Spainisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Collaboration in software developmentccs: Social and professional topics Socio-technical systems
@@ 1. Introduction
Software practitioners encounter a range of problems at work. Some of these problems relate to their work, for example, incomplete requirements, software testing, and code-related problems (Ragkhitwetsagul et al., 2022). Other problems relate to human and social aspects, for example, personality clashes inhibiting collaboration (Fernando Capretz, 2014) and the need for improved customer relations and communications (Lenberg et al., 2015). One such factor is fairness at work. Studies show that understanding fairness problems can improve software products that can indirectly impact software companies’ profit (Fernando Capretz, 2014).
German et al. (German et al., 2018) developed a framework to study the perceptions of fairness in modern code reviews. Another study found differences between male and female practitioners when expressing sentiments during Software Engineering (SE) interactions and how their opinions differ based on the gender of the recipient indicating gender bias (Paul et al., 2019). Further, there are studies investigating the challenges women face in global software development teams (Trinkenreich et al., 2022) and in open source software development (Imtiaz et al., 2019).
While previous studies (German et al., 2018; Paul et al., 2019; Trinkenreich et al., 2022; Imtiaz et al., 2019; Gunawardena et al., 2022) have focused on specific issues related to fairness in SE, there remains a broader gap in the literature. Our study explores the overall concept of fairness and understanding the specific fairness-related concerns of software practitioners. We intend to align our understanding of fairness relating to human and social factors in SE with the established principles of Organizational Justice (OJ) literature (Colquitt, 2001), a framework that has previously informed fairness exploration in SE work by German et al. (German et al., 2018). Fairness refers to the perception that the practices, policies, and interactions within an organization are unbiased, and just (Colquitt, 2001). Fairness relating to human and social factors in SE encompass, for example, fair task allocation, respectful interactions, and clear communication within engineering teams to promote a fair and supportive work environment.
@@
RQ1: How often is fairness, relating to human and social factors in software engineering, discussed on Stack Exchange sites? We analyzed eight StEx sites likely to discuss fairness in human and social factors in SE. This includes Stack Overflow, Software Engineering StEx, Workplace StEx, Code Review StEx, Open Source StEx, Computer Science StEx, DevOps StEx, and Project Management StEx. We identified 1,354 candidate fairness posts across five StEx sites, of which 136 posts (10.0%) relate to fairness in human and social factors in SE. This semi-automated process employs a mixed-methods empirical approach, consisting of keyword searches and manual analysis. RQ1 reveals that the StEx sites have various discussions on fairness, covering areas such as algorithms, computation, and human and social factors. We find that 10 out of every 100 candidate fairness posts relate to human and social aspects of SE, with most of them being hosted on Workplace StEx.
The second research question explores the fairness problems discussed by software practitioners and their characteristics.
RQ2: What characterizes fairness discussions on Stack Exchange sites? We characterize fairness discussions on three attributes:
RQ2.1: What are the contexts in which fairness is discussed? We manually identified 10 specific fairness contexts and categorized the fairness posts accordingly. The results indicate that a majority of posts (24 out of 136 posts) discuss fairness in income. Then, we map the selected posts to the fairness dimensions borrowed from OJ literature:
RQ2.2: Which dimensions of fairness are discussed? To this end, we employed Colquitt’s fairness measure (Colquitt, 2001), widely used in OJ literature (German et al., 2018). Our study showed that practitioners are most concerned about fairness in decision-making processes influencing the outcome (53 fairness posts or 38.97%), consistent with what has been observed generally in OJ literature. Last, we investigate the demographics of the post owners for characterization in order to understand the potential relationship between user characteristics and the fairness posts:
@@ 2.1. Mining Stack Exchange Posts
Stack Exchange (StEx) is a network of 181 communities that are created and run by experts and enthusiasts who are passionate about a specific topic (Exchange, 2022c). Stack Overflow (SO), an online platform found in 2008 where programmers could help answer each other problems for free, was the origin of StEx. Ever since, the network has expanded to include many other communities, such as technology (e.g., Game Development), professional (e.g., Workplace), and business (e.g., Quantitative Finance).
On StEx sites, a user can ask questions and receive answers from other users. Good answers are voted up and shown first, making them easy to locate. The owner of the question can mark one answer as ”accepted,” indicating that it worked for the person who asked. Comments can be used to request additional information or to clarify a question or answer.
@@ 2.2. Fairness in software engineering
Several studies investigate how bias and unfairness prevail. Trinkenreich et al. (Trinkenreich et al., 2022) surveyed Ericsson software developers and found that socio-cultural barriers prevent women from advancing in global software development teams. Survey participants indicate that women’s voices are silenced by men and they are included in projects based on gender rather than ability (Trinkenreich et al., 2022). In code review, Paul et al. (Paul et al., 2019) found that male developers in three out of six projects were writing more negative comments and refusing to assist their female peers, leading to disparities in sentiment expressions between male and female practitioners. Similarly, Gunawardena et al. (Gunawardena et al., 2022) discovered that destructive critique in code reviews affects participants’ moods, desire to continue working, and written reactions to the criticism. Women are found to perceive destructive criticism as less acceptable and are less motivated to work with the developer after receiving it. Therefore, destructive criticism in code reviews may contribute to the lack of gender diversity observed in the software industry.
@@ 3. Methodology
Step 1: Identify sites. The question we aimed to answer in this study is: what are the fairness problems that software practitioners face during software development? Gaining insights into these problems can be difficult since they often involve sensitive information based on personal experiences. StEx sites, which host discussions on a wide range of topics, offer valuable data, including conversations about fairness in SE. We believe StEx’s long standing role in the development community makes it an ideal starting point for our study. The anonymity provided to users encourages candid discussions about their experiences (Alnasser et al., 2021) making the information a direct reflection of the experiences faced by users.
@@
Keyword Search. We compiled a list of keywords and phrases likely related to concerns about fairness, such as bias, justice, equality, and equity. Inspired by similar studies (Krishtul et al., 2022; Obie et al., 2020), the keywords were also chosen based on our observations from the 74 posts, as well as a list of Colquitt’s fairness rules (e.g., equity, respect, bias suppression) (Colquitt, 2001; Colquitt and Chertkoff, 2002) and their associated synonyms and antonyms. We used entries from the online Thesaurus (the, 2023) and Merriam-Webster (mer, 2023) dictionary to identify the associated synonyms and antonyms.
To improve the search and deal with coherency between words, we only selected semantically and contextually linked words. For example, the word right is synonym for equity. However, the two words are not always used in the same context or with the same intended meaning, which can increase false positive rate. Therefore, we omitted right in our keywords list. The second query yielded 121,560 posts in total, excluding the posts found in first extraction.
@@ 4. Findings
RQ1: How often is fairness, relating to human and social factors in software engineering, discussed on Stack Exchange sites?
@@
RQ2: What characterizes fairness discussions on Stack Exchange sites?
@@ 6. Threats to Validity
Construct Validity refers to how well a measurement captures its intended aspects (Ralph and Tempero, 2018). However, applying Colquitt’s fairness measure in RQ2.2 could be influenced by researchers bias. To mitigate this threat, we conducted consensus coding process between 3 authors, and consulted seminal papers in justice literature (Colquitt, 2001; Colquitt and Chertkoff, 2002; Colquitt, 2012; Colquitt et al., 2013; Colquitt and Rodell, 2015) as well as previous research that applied fairness theory in SE (German et al., 2018). The results of the study was confirmed by OJ researcher, who agreed that procedural fairness is the most crucial fairness dimension, which supports our observation in RQ2.2.
@@
Internal Validity is described as threats that may have influenced our findings (Wohlin et al., 2012). We randomly selected 300 posts from each of the eight StEx sites (1,860 posts total) during the second post filtering, manually extracted fairness posts, and performed a qualitative analysis for RQ2.1. However, our findings may be threatened by several factors. We may have missed crucial posts about fairness in SE due to our random selection of 1,860 posts. The manual extraction of fairness posts may also be subjective and prone to mistakes, although we took steps to reduce this threat by having multiple authors involved in the process. We addressed the potential issue of low numbers of selected posts by filtering and manually extracting the data twice, which improved the reliability of our selection. This careful process not only enhanced our current analysis but also laid the groundwork for creating an automated solution in the future. Additionally, the process of categorizing posts into contexts in RQ2.1 may be prone to mistakes, but we validated our work by having three authors hold several discussions before finalizing the list of contexts. Lastly, we acknowledge the limitations in identifying fairness posts from 8 StEx sites. However, we concentrated our data collection efforts on sites that we believed were most likely to provide valuable insights into our research questions. We also made attempts to uncover the possible explanations for why we found that three of the sites did not have any fairness posts on Section 5.1.
@@
External Validity is the degree to which our findings can be applied to different situations (Wohlin et al., 2012). We analyzed 136 fairness posts on eight StEx sites. Our findings are based only on these posts and may not apply to all posts on the eight StEx sites nor to overall SE concerns. Additionally, they cannot be generalized to other grey literatures. However, the finding that procedural and distributive fairness are the two dimensions that practitioners are most concerned about may apply more broadly because they reflect what prior studies have reported (German et al., 2018). In RQ2.1, the interpretation of context depends on the post, and one way to diversify information sources is to consider several StEx sites. Therefore, the list of contexts identified in this study does not necessarily reflect the opinions of software practitioners beyond the current sample. This study also did not identify the opinions of practitioners from minoritized groups who might have had different experiences related to fairness.
@@ 7. Conclusion and future work
We conducted a study on fairness in SE by analyzing StEx posts. We examined 4,178 posts from eight StEx sites to find discussions about fairness related to social and human factors. We further explored the contexts where the problem arises and the post owners’ characteristics. Using Colquitt’s list of questions, we also studied the fairness dimensions that software practitioners care about.
Our findings reveal that 136 posts addressed fairness concerns in processes and interactions related to SE, offering valuable insights into SE practice. The diverse range of posts analyzed has uncovered unique fairness perspectives. By focusing on the qualitative data, we have gained a deeper understanding of the fairness problems experienced by several practitioners in the SE field. In conclusion, this study highlights the significance of considering diverse perspectives and understanding fairness issues in SE practice. By addressing these issues, we can strive towards a more fair SE community. In the future, we intend to expand our exploratory study to observe the solution to the fairness problems and how fairness affects the outcome of SE.

### v3

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 18th ACM / IEEE International Symposium on Empirical Software Engineering and Measurement; October 24–25, 2024; Barcelona, Spainbooktitle: Proceedings of the 18th ACM / IEEE International Symposium on Empirical Software Engineering and Measurement (ESEM ’24), October 24–25, 2024, Barcelona, Spaindoi: 10.1145/3674805.3686687isbn: 979-8-4007-1047-6/24/10ccs: Software and its engineering Collaboration in software developmentccs: Social and professional topics Socio-technical systems
@@ 1. Introduction
Software practitioners encounter a range of problems at work. Some of these problems relate to their work, for example, incomplete requirements, software testing, and code-related problems (Ragkhitwetsagul et al., 2022). Other problems relate to human and social aspects, for example, personality clashes inhibiting collaboration (Fernando Capretz, 2014) and the need for improved customer relations and communications (Lenberg et al., 2015). One such factor is fairness at work. Studies show that understanding fairness problems can improve software products that can indirectly impact software companies’ profit (Fernando Capretz, 2014). German et al. (German et al., 2018) developed a framework to study the perceptions of fairness in modern code reviews. Another study found differences between male and female practitioners when expressing sentiments during Software Engineering (SE) interactions and how their opinions differ based on the gender of the recipient indicating gender bias (Paul et al., 2019). Further, there are studies investigating the challenges women face in global software development teams (Trinkenreich et al., 2022b) and in open source software development (Imtiaz et al., 2019).
While previous studies (German et al., 2018; Paul et al., 2019; Trinkenreich et al., 2022b; Imtiaz et al., 2019; Gunawardena et al., 2022) have focused on specific issues related to fairness in SE, there remains a broader gap in the literature. These studies highlight the potential for improving developer satisfaction and productivity through more fair work practices. Our study aims to explore the overall concept of fairness and understand the specific fairness-related concerns of software practitioners, providing a comprehensive view that builds on these previous insights. We align our understanding of fairness relating to human and social factors in SE by building upon the theoretical framework provided by Organizational Justice (OJ) literature (Colquitt, 2001), a framework that has previously informed fairness exploration in SE work by German et al. (German et al., 2018). Fairness refers to the perception that the practices, policies, and interactions within an organization are unbiased, and just (Colquitt, 2001). Fairness relating to human and social factors in SE encompass, for example, fair task allocation, respectful interactions, and clear communication within engineering teams to promote a fair and supportive work environment.
@@
RQ1: How often is fairness, relating to human and social factors in SE, discussed on StEx sites? We analyzed eight StEx sites likely to discuss fairness in human and social factors in SE. This includes Stack Overflow, Software Engineering StEx, Workplace StEx, Code Review StEx, Open Source StEx, Computer Science StEx, DevOps StEx, and Project Management StEx. We identified 1,354 candidate fairness posts across five StEx sites, of which 136 posts (10.0%) relate to fairness in human and social factors in SE. This semi-automated process employs a mixed-methods empirical approach, consisting of keyword searches and manual analysis. RQ1 reveals that the StEx sites have various discussions on fairness, covering areas such as algorithms, computation, and human and social factors. We find that 10 out of every 100 candidate fairness posts relate to human and social aspects of SE, with most of them being hosted on Workplace StEx. The second research question explores the fairness problems discussed by software practitioners and their characteristics.
RQ2: What characterizes fairness discussions on StEx sites? We characterize fairness discussions on three attributes.
RQ2.1: What are the contexts in which fairness is discussed? We manually identified 10 specific fairness contexts and categorized the fairness posts accordingly. The results indicate that a majority of posts (24 out of 136 posts) discuss fairness in income. Then, we map the selected posts to the fairness dimensions borrowed from OJ literature.
RQ2.2: Which dimensions of fairness are discussed? To this end, we employed Colquitt’s fairness measure (Colquitt, 2001), widely used in OJ literature (German et al., 2018). Our study showed that practitioners are most concerned about fairness in decision-making processes influencing the outcome (53 fairness posts or 38.97%), consistent with what has been observed generally in OJ literature. Last, we investigate the demographics of the post owners for characterization in order to understand the potential relationship between user characteristics and the fairness posts.
@@ 2.1. Mining Stack Exchange Posts
Stack Exchange (StEx) is a network of 181 communities that are created and run by experts and enthusiasts who are passionate about a specific topic (Exchange, 2022c). Stack Overflow (SO), an online platform found in 2008 where programmers could help answer each other problems for free, was the origin of StEx. Ever since, the network has expanded to include many other communities, such as technology (e.g., Game Development), professional (e.g., Workplace), and business (e.g., Quantitative Finance). On StEx sites, a user can ask questions and receive answers from other users. Good answers are voted up and shown first, making them easy to locate. The owner of the question can mark one answer as ”accepted,” indicating that it worked for the person who asked. Comments can be used to request additional information or to clarify a question or answer.
@@ 2.2. Fairness in Software Engineering
Several studies investigate how bias and unfairness prevail. Trinkenreich et al. (Trinkenreich et al., 2022a) surveyed Ericsson software developers and found that socio-cultural barriers prevent women from advancing in global software development teams. Survey participants indicate that women’s voices are silenced by men and they are included in projects based on gender rather than ability (Trinkenreich et al., 2022a). In code review, Paul et al. (Paul et al., 2019) found that male developers in three out of six projects were writing more negative comments and refusing to assist their female peers, leading to disparities in sentiment expressions between male and female practitioners. Similarly, Gunawardena et al. (Gunawardena et al., 2022) discovered that destructive critique in code reviews affects participants’ moods, desire to continue working, and written reactions to the criticism. Women are found to perceive destructive criticism as less acceptable and are less motivated to work with the developer after receiving it. Therefore, destructive criticism in code reviews may contribute to the lack of gender diversity observed in the software industry.
@@ 3. Methodology
Step 1: Identify sites. The goal in this study is to understand the fairness problems that software practitioners face during software development. Gaining insights into these problems can be difficult since they often involve sensitive information based on personal experiences. StEx sites, which host discussions on a wide range of topics, offer valuable data, including conversations about fairness in SE. We believe StEx’s long standing role in the development community makes it an ideal starting point for our study. The anonymity provided to users encourages candid discussions about their experiences (Alnasser et al., 2021) making the information a direct reflection of the experiences faced by users.
@@
Keyword Search. We compiled a list of keywords and phrases likely related to concerns about fairness, such as bias, justice, equality, and equity. Inspired by similar studies (Krishtul et al., 2022; Obie et al., 2020), the keywords were also chosen based on our observations from the 74 posts, as well as a list of Colquitt’s fairness rules (e.g., equity, respect, bias suppression) (Colquitt, 2001; Colquitt and Chertkoff, 2002) and their associated synonyms and antonyms. We used entries from the online Thesaurus (the, 2023) and Merriam-Webster (mer, 2023) dictionary to identify the associated synonyms and antonyms. To improve the search and deal with coherency between words, we only selected semantically and contextually linked words. For example, the word right is synonym for equity. However, the two words are not always used in the same context or with the same intended meaning, which can increase false positive rate. Therefore, we omitted right in our keywords list. The second query yielded 121,560 posts in total, excluding the posts found in first extraction.
@@ 4. Findings
RQ1: How often is fairness, relating to human and social factors in SE, discussed on StEx sites?
@@
RQ2: What characterizes fairness discussions on StEx sites?
@@ 6. Threats to Validity
Construct Validity refers to how well a measurement captures its intended aspects (Ralph and Tempero, 2018). However, applying Colquitt’s fairness measure in RQ2.2 could be influenced by researchers bias. To mitigate this threat, we conducted consensus coding process between 3 authors, and consulted seminal papers in OJ literature (Colquitt, 2001; Colquitt and Chertkoff, 2002; Colquitt, 2012; Colquitt et al., 2013; Colquitt and Rodell, 2015) as well as previous research that applied fairness theory in SE (German et al., 2018). The results of the study was confirmed by OJ researcher, who agreed that procedural fairness is the most crucial fairness dimension, which supports our observation in RQ2.2.
@@
Internal Validity is described as threats that may have influenced our findings (Wohlin et al., 2012). We randomly selected 300 posts from each of the 8 StEx sites (1,860 posts total) during the second post filtering, manually extracted fairness posts, and performed a qualitative analysis for RQ2.1. However, our findings may be threatened by several factors. We may have missed crucial posts about fairness in SE due to our random selection of 1,860 posts. The manual extraction of fairness posts may also be subjective and prone to mistakes, although we took steps to reduce this threat by having multiple authors involved in the process. We addressed the potential issue of low numbers of selected posts by filtering and manually extracting the data twice, which improved the reliability of our selection. This careful process not only enhanced our current analysis but also laid the groundwork for creating an automated solution in the future. Additionally, the process of categorizing posts into contexts in RQ2.1 may be prone to mistakes, but we validated our work by having 3 authors hold several discussions before finalizing the list of contexts. Lastly, we acknowledge the limitations in identifying fairness posts from 8 StEx sites. However, we concentrated our data collection efforts on sites that we believed were most likely to provide valuable insights into our research questions. We also made attempts to uncover the possible explanations for why we found that 3 of the sites did not have any fairness posts on Section 5.1.
@@
External Validity is the degree to which our findings can be applied to different situations (Wohlin et al., 2012). We analyzed 136 fairness posts on 8 StEx sites. Our findings are based only on these posts and may not apply to all posts on the 8 StEx sites nor to overall SE concerns. Additionally, they cannot be generalized to other grey literatures. However, the finding that procedural and distributive fairness are the two dimensions that practitioners are most concerned about may apply more broadly because they reflect what prior studies have reported (German et al., 2018). In RQ2.1, the interpretation of context depends on the post, and one way to diversify information sources is to consider several StEx sites. Therefore, the list of contexts identified does not necessarily reflect the opinions of software practitioners beyond the current sample. This study also did not identify the opinions of practitioners from minoritized groups who might have had different experiences related to fairness.
@@ 7. Conclusion and future work
We conducted a study on fairness in SE by analyzing StEx posts. We examined 4,178 posts from eight StEx sites to find discussions about fairness related to social and human factors. We further explored the contexts where the problem arises and the post owners’ characteristics. Using Colquitt’s list of questions, we also studied the fairness dimensions that practitioners care about.
Our findings reveal that 136 posts addressed fairness concerns in processes and interactions related to SE.The diverse range of posts analyzed has uncovered unique fairness perspectives. By focusing on the qualitative data, we have gained a deeper understanding of the fairness problems experienced by several practitioners in the SE field. In conclusion, this study highlights the significance of considering diverse perspectives and understanding fairness issues in SE practice. By addressing these issues, we can strive towards a more fair SE community. In the future, we intend to expand our exploratory study to observe the solution to the fairness problems and how fairness affects the outcome of SE.